# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v4

@@ 5. Post-Optimization and Usage
The NLP determines the trip count for each loop, array partitioning, and on-chip buffer sizes. Functions for loading/storing data between off-chip and on-chip memory are automatically generated with the maximum burst size. Further optimizations are also applied to improve QoR. Optimization for Non-Constant Trip Count  To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler. We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n. d.]). Next, we replace all statements iterated by this loop with calls to a function. For example, instead of using $C[i][i]+=B[i][i]$, we replace it with $C[i][j]=f(C[i][j],B[i][i])$, where the function $f$ ensures compliance with the constraints of the non-constant trip count loop. This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the output value. Therefore, we employ these techniques to reduce compilation time and achieve designs with a good QoR. Grouping Data Transfers  In the NLP, we already account for parallel transfers of arrays when they are transferred at the same level in the code (e.g., within the same loop). However, we do not consider parallel transfers between arrays belonging to different loop bodies. In the 2mm example, we first handle the transfer of arrays required for the initial matrix multiplication, only transferring the arrays for the second multiplication after the first is complete. However, FPGA DRAM banks enable parallel transfers of different arrays. Therefore, we can optimize load and store operations during post-processing to enhance memory transfer overlap. Code Transformation for HLS  To simplify the understanding of the HLS compiler, particularly to ensure that the compiler can pipeline loops with the correct initiation interval (II), we simplify the reductions (regardless of whether we use tree reductions) in the fine-grained unrolled part (level 2). Therefore, under the pipeline we create a variable to accumulate the reduction, which we then add to the output. Additionally, we use the loop_flatten off pragma on all reduction loops above the pipeline. Although this adjustment could theoretically improve the QoR, in practice, flattening these loops can complicate the code and potentially increase the initiation interval (II) of the pipeline. Usage  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources. To fully benefit developers, our NLP formulation allows the template to be completely customized to the user’s needs. Each variable in the design space can be defined by the user if they have specific constraints. For example, in the gemm template (Listing 3), the user can set the permutations for level 0 while allowing the NLP to explore other parameters. Additionally, they can specify the on-chip size of matrix A, enabling exploration within the defined constraints.
@@ 6. Evaluation
The goal of this evaluation is to demonstrate how effectively our approach applies code transformations and inserts hardware directives while respecting user-defined resource constraints. We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2021), NLP-DSE (Pouget et al., 2024b), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022). Since generating a bitstream without significant manual intervention is challenging for some frameworks, we first demonstrate our ability to achieve a high-quality design by utilizing all available resources—ensuring fair comparison with the other frameworks—without considering place and route. We generate HLS reports using the vitis-flow, which accounts for memory transfers between off-chip and on-chip, necessitating the inclusion of memory transfer steps in ScaleHLS’s code. Thus, we show that our designs are synthesizable by targeting the resources of a single SLR, which eliminates the need for manual partitioning. Due to the time required for bitstream generation, we restrict our evaluation to a subset of the kernels.
@@ 6.1. Setup
We conducted experiments using kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]), a CNN kernel, and matrix multiplication tasks similar to those in the BERT transformer model. All computations used single-precision floating-point data for direct comparison with frameworks like AutoDSE, NLP-DSE, and ScaleHLS, using medium-sized Polybench/C datasets. We also compare our approach with HARP (Sohrabizadeh et al., 2023), which enhances HLS design exploration using a Graph Neural Network (GNN) model to predict HLS tool behavior. HARP explores this space for one hour and synthesizes the top 10 designs. We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions. As detailed in (Pouget et al., 2024b), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes. We selected these problem sizes to highlight how our method performs in transforming code and inserting pragmas in cases where full unrolling is not feasible. We also demonstrate the effectiveness of our framework in tiling large tasks such as CNN and bert_3072_100, ensuring that only a small portion (a tile) of each array fits on-chip, in line with memory constraints. For the CNN kernel, the problem size and loop dimensions are the same as described in Section 2. For bert_n_m matrix multiplication, the dimensions are I = n, J = m, K = n (reduction). We used the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to demonstrate our method’s effectiveness, generating reports with Vitis 2023.2. When enabling tree reduction, we selected the “funsafe math optimizations” option to allow commutative and associative reduction operators for logarithmic time reductions. For our hardware setup, we targeted the Xilinx Alveo U200 device running at 250 MHz. We use the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL. The experiments were performed on an AMD EPYC 7V13 64-core processor, with 32 GB of memory and 4 cores allocated for each run.
@@ 6.2. Experimental Evaluation
To ensure a fair comparison, we modified kernels with non-constant trip counts for all benchmarks, as detailed in Section 5, except for ScaleHLS, which already handles such loops. We generated AutoDSE’s design space using the ds_generator command, applying all trip count factors for unroll and tile sizes. Since AutoDSE lacks array partitioning, its design space is slightly larger. AutoDSE was run with a 1,000-minute exploration timeout and a 180-minute HLS synthesis timeout. For NLP-DSE, we followed the original paper’s parameters. HARP was run with the authors’ settings using Vitis 2021.1 and tree reduction enabled. ScaleHLS assumes that all kernel data are already located in on-chip memory. To provide a fair comparison, we modified their code to handle off-chip to on-chip memory transfers. We optimized these transfers to 512 bits per cycle, with all arrays transferred in parallel. Each array was assigned to a different DRAM bank. However, due to the extensive manual modifications required, we did not extend this comparison to all kernels. For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds (4 hours) per design. By default, we assume optimistic resource reuse between non-parallel statements. If DSP utilization exceeds limits, we rerun the NLP with pessimistic assumptions (no resource reuse) and regenerate the design. The maximum partitioning constraint, $max_{part}$, was set to 1024, per AMD/Xilinx guidelines. Each synthesis starts with a C-simulation to verify correctness. We generate Vitis reports using the vitis-flow, which counts memory transfers for all frameworks and considers the resources of the entire FPGA. For bitstream generation, we consider only one SLR to eliminate any manual intervention. We consider 60% of the resources of a single SLR to quickly identify a synthesizable design, minimizing the number of times AutoDSE needs to be run.
@@ 6.3.1. AutoDSE, NLP-DSE and Scale-HLS with tree reduction
Table 3 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE), Scale-HLS (S-HLS) and Sisyphus (ours) with tree reduction. The T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation. The column Perf. Impr. of ours framework vs. displays the performance improvement of our framework compared to the three different frameworks. Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%. For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq. 17), as the constraint with optimistic reuse (Eq. 16) resulted in kernels with resource over-utilization. For Gesummv, ScaleHLS is 1.41x faster. However, 91% of our design’s execution time is spent on data transfers. ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing. Our framework uses 64-bit chunks, the largest bitwidth that divides the array. With 512-bit transfers, our framework could achieve 14.50 GF/s. For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers.
@@
Table 3. Throughput (GF/s) Comparison of Sisyphus, AutoDSE, NLP-DSE, and ScaleHLS with Tree Reduction Using Vitis HLS
Sisyphus averages a speedup of 683.91x, 641.46x and 434.78x over AutoDSE, NLP-DSE and Scale-HLS respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 9.17x, 3.47x and 16.15x. Preserve the original schedule when necessary  For atax, Sisyphus preserves the original schedule and inserts pragmas, effectively maintaining the existing schedule while efficiently incorporating pragmas, similar to two DSEs focused on pragma insertion. Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit. Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels. For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities. Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, syr2k, and trmm, the original schedule imposes constraints on the achievable level of parallelism. This limitation highlights the performance enhancements made possible by Sisyphus. We observe QoR improvements for these kernels because the code transformations enable a combination of loop unrolling that cannot be achieved even with maximal loop distribution, as each loop can be partially unrolled.
@@ 6.3.2. AutoDSE and NLP-DSE without tree reduction
Transformations without tree reduction, which may be needed to preserve floating-point data precision, must be different since the optimizations differ when tree reductions are not involved. Table 4 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours) without tree reduction. Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.53x over NLP-DSE across the evaluated kernels without tree reduction. The geometric mean speedup is 18.50x and 4.80x, respectively. On average, the performance improvement without tree reduction is greater than with tree reduction, particularly for gemm (Listing 2) and gemver. This is because, without tree reduction, optimizing the reduction loop becomes challenging, making code transformations even more necessary. However the difference observed between syrk and syr2k with tree reduction is due to the HLS compiler achieving an II=2 instead of the expected II=1. However, without tree reduction, the compiler successfully achieves II=1.
@@
Average 133.84 676.03x 567.53x
Geo Mean 41.62 18.50x 4.80x
Table 4. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE without Tree Reduction Using Vitis HLS
@@ 6.3.3. AutoDSE and NLP-DSE for Bert and CNN
Table 5 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours). The TR column denotes the status of the tree reduction option. For bert_100_64, we observed a minor decrease in QoR, attributed to the NLP selecting a configuration where the pipeline was not applied as expected. Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred. In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
@@
Table 5. Throughput (GF/s) Comparison of Sisyphus, AutoDSE and NLP-DSE Using Vitis HLS
@@ 6.3.4. HARP
In Table 6 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023). The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE. Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin source-to-source compiler, which facilitates loop permutations when partially unrolling consecutive loops. This allows them to find the same schedule that our framework identified. However, for Gemm, the design discovered by HARP falls outside our specified parameter space, particularly because of our restriction on maximal array partitioning. Notably, it is the only design in the top 10 that outperforms our framework.
@@
bicg 0.92 1.96 75,25,35,36 38,65,53,34 2.13x
gemm 125.59 88.61 29,80,55,40 32,48,24,17 0.71x
gemver 1.66 9.42 29,28,35,19 30,64,65,33 5.69x
mvt 7.07 8.24 40,78,43,30 53,43,49,30 1.17x
Average 27.39 22.04 2.16x
Geo Mean 4.71 7.66 1.62x

### v5

@@ 5. Post-Optimization and Customization
The NLP determines the trip counts, array partitioning, and on-chip buffer sizes. Data transfer functions between off-chip and on-chip memory are automatically generated with the maximum burst size. Further optimizations are also applied to improve QoR. Optimization for Non-Constant Trip Count  To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler. We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n. d.]). Next, we replace all statements iterated by this loop with calls to a function. For example, instead of using $C[i][i]+=B[i][i]$, we replace it with $C[i][j]=f(C[i][j],B[i][i])$, where the function $f$ ensures compliance with the constraints of the non-constant trip count loop. This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the original output value. This technique helps to reduce compilation time and achieve designs with a good QoR. Grouping Data Transfers  In the NLP, we already account for parallel transfers of arrays when they are transferred at the same level in the code (e.g., within the same loop). However, we do not consider parallel transfers between arrays belonging to different loop bodies. In our example, we first manage the transfer of arrays required for initializing the matrix, transferring the arrays for multiplication only after the initialization is complete. However, FPGA DRAM banks allow parallel transfers of different arrays. Consequently, we can optimize load and store operations during post-processing to improve memory transfer overlap. Code Transformation for HLS  To help the HLS compiler pipeline loops with the correct initiation interval (II), we simplify reductions at level 2. A variable accumulates the reduction within the pipeline, which is then added to the output. Additionally, we use the loop_flatten off pragma on all reduction loops above the pipeline. Although this adjustment could theoretically improve the QoR, in practice, flattening these loops can complicate the code and potentially increase the initiation interval (II) of the pipeline. Customization  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources. To fully benefit developers, our NLP formulation allows the template to be completely customized to the user’s needs. Each variable in the design space can be defined by the user if they have specific constraints. For example, in the gemm template (Listing 3), the user can set the permutations for level 0 while allowing the NLP to explore other parameters. Additionally, they can specify the on-chip size of matrix $A$ by directly defining a variable in the NLP, allowing for exploration within the defined constraints.
@@ 6. Evaluation
The goal of this evaluation is to demonstrate how effectively our approach applies code transformations and inserts hardware directives while respecting user-defined resource constraints. We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2022b), NLP-DSE (Pouget et al., 2025), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022). Since generating a bitstream without significant manual intervention is challenging for some frameworks, we first demonstrate our ability to achieve a high-quality design by utilizing all available resources—ensuring fair comparison with the other frameworks—without considering place and route. We generate HLS reports using the vitis-flow, which accounts for memory transfers between off-chip and on-chip, necessitating the inclusion of memory transfer steps in ScaleHLS’s code. Thus, we show that our designs are synthesizable by targeting the resources of a single SLR, which eliminates the need for manual partitioning. Due to the time required for bitstream generation, we restrict our evaluation to a subset of the kernels.
@@ 6.1. Setup
We conducted experiments using kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]), a CNN kernel, and matrix multiplication tasks similar to those in the BERT transformer model. All computations used single-precision floating-point data for direct comparison with frameworks like AutoDSE, NLP-DSE, and ScaleHLS, using medium-sized Polybench/C datasets. We also compare our approach with HARP (Sohrabizadeh et al., 2023), which enhances HLS design exploration using a Graph Neural Network (GNN) model to predict HLS tool behavior. HARP explores this space for one hour and synthesizes the top 10 designs. We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions. As detailed in (Pouget et al., 2025), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes. We selected these problem sizes to highlight how our method performs in transforming code and inserting pragmas in cases where full unrolling is not feasible. We also demonstrate the effectiveness of our framework in tiling large tasks such as CNN and bert_3072_100, ensuring that only a small portion (a tile) of each array fits on-chip, in line with memory constraints. For the CNN kernel, the problem size and loop dimensions are the same as described in Section 2. For bert_n_m matrix multiplication, the dimensions are I = n, J = m, K = n (reduction). We used the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to demonstrate our method’s effectiveness, generating reports with Vitis 2023.2. When enabling tree reduction, we selected the “funsafe math optimizations” option to allow commutative and associative reduction operators for logarithmic time reductions. For our hardware setup, we targeted the Xilinx Alveo U200 device operating at 250 MHz. We utilized the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL.
@@ 6.2. Experimental Evaluation
To ensure a fair comparison, we modified kernels with non-constant trip counts for all benchmarks, as detailed in Section 5, except for ScaleHLS, which already handles such loops. We generated AutoDSE’s design space using the ds_generator command, applying all trip count factors for unroll and tile sizes. Since AutoDSE lacks array partitioning, its design space is slightly larger. AutoDSE was run with a 1,000-minute exploration timeout and a 180-minute HLS synthesis timeout. For NLP-DSE, we followed the original paper’s parameters (Pouget et al., 2025). HARP was run with the authors’ settings (Sohrabizadeh et al., 2023) using Vitis 2021.1 and tree reduction enabled. ScaleHLS assumes that all kernel data are already located in on-chip memory. To provide a fair comparison, we modified their code to handle off-chip to on-chip memory transfers. We optimized these transfers to 512 bits per cycle, with all arrays transferred in parallel. Each array was assigned to a different DRAM bank. However, due to the extensive manual modifications required, we did not extend this comparison to all kernels. For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds per design. By default, we assume optimistic resource reuse between non-parallel statements. If DSP utilization exceeds limits, we rerun the NLP with pessimistic assumptions (no resource reuse) and regenerate the design. The maximum partitioning constraint, $max_{part}$, was set to 1024, per AMD/Xilinx guidelines. Each synthesis starts with a C-simulation to verify correctness. We generate Vitis reports using the vitis-flow, which counts memory transfers for all frameworks and considers the resources of the entire FPGA. For bitstream generation, we consider only one SLR to eliminate any manual intervention. We consider 60% of the resources of a single SLR to quickly identify a synthesizable design, minimizing the number of times AutoDSE needs to be run.
@@ 6.3.1. AutoDSE, NLP-DSE and Scale-HLS with tree reduction
Table 3 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE), Scale-HLS (S-HLS) and Sisyphus (ours) with tree reduction. The T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation. The column Perf. Impr. of ours framework vs. displays the performance improvement of our framework compared to the three different frameworks.
@@
Table 3. Throughput (GF/s) Comparison with Tree Reduction Using Vitis HLS
Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%. For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq. 17), as the constraint with optimistic reuse (Eq. 16) resulted in kernels with resource over-utilization. For Gesummv, ScaleHLS is 1.41x faster. However, 91% of our design’s execution time is spent on data transfers. ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing. Our framework uses 64-bit chunks, the largest bitwidth that divides the array. With 512-bit transfers, our framework could achieve 14.50 GF/s. For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers. Sisyphus averages a speedup of 683.91x, 641.46x and 434.78x over AutoDSE, NLP-DSE and Scale-HLS respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 9.17x, 3.47x and 16.15x. Preserve the original schedule when necessary  For atax, Sisyphus preserves the original schedule and inserts pragmas, effectively maintaining the existing schedule while efficiently incorporating pragmas, similar to two DSEs focused on pragma insertion. Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit. Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels. For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities. Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, syr2k, and trmm, the original schedule imposes constraints on the achievable level of parallelism. This limitation highlights the performance enhancements made possible by Sisyphus. We observe QoR improvements for these kernels because the code transformations enable a combination of loop unrolling that cannot be achieved even with maximal loop distribution, as each loop can be partially unrolled.
@@ 6.3.2. AutoDSE and NLP-DSE without tree reduction
Transformations without tree reduction, which may be needed to preserve floating-point data precision, must be different since the optimizations differ when tree reductions are not involved. Table 4 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours) without tree reduction.
@@
Average 133.84 676.03x 567.10x
Geo Mean 41.62 18.50x 4.49x
Table 4. Throughput (GF/s) Comparison without Tree Reduction Using Vitis HLS
Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.10x over NLP-DSE across kernels without tree reduction, with geometric means of 18.50x and 4.49x, respectively. Performance improvement is higher without tree reduction, especially for gemm (Listing 2) and gemver, as optimizing reduction loops becomes harder, increasing the need for code transformations. The difference between syrk and syr2k with tree reduction occurs because the HLS compiler achieves II=2 instead of the expected II=1, while without tree reduction, II=1 is achieved.
@@ 6.3.3. AutoDSE and NLP-DSE for Bert and CNN
Table 5 compares AutoDSE (A-DSE), NLP-DSE (N-DSE), and Sisyphus (ours). The TR column shows the tree reduction status. For bert_100_64, we observed a minor decrease in QoR, attributed to the NLP selecting a configuration where the pipeline was not applied as expected.
@@
Table 5. Throughput (GF/s) Comparison Using Vitis HLS
Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred. In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
@@ 6.3.4. HARP
In Table 6 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023).
@@
bicg 0.92 1.96 75,25,35,36 38,65,48,31 2.13x
gemm 125.59 87.86 29,80,55,40 32,48,36,23 0.70x
gemver 1.66 9.42 29,28,35,19 34,8,31,11 5.69x
mvt 7.07 8.29 40,78,43,30 53,44,57,30 1.17x
Average 27.39 20.71 1.45x
Geo Mean 4.71 6.28 1.33x
@@
The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE. Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin compiler, which facilitates loop permutations when partially unrolling consecutive loops. This allows them to find the same schedule that our framework identified. However, for gemm, the design discovered by HARP falls outside our specified parameter space, due to our restriction on maximum array partitioning.